What mass (in grams) of CuCl2 would be needed to make 575 mL of 1.12 M solution?


1
Expert's answer
2021-03-22T06:21:25-0400

Molar Mass Cu = 63.5 ×1 = 63.5

Molar Mass Cl = 35.5 × 2 = 71

Molar Mass CuCl2 = 134.5

1.12 × 134.5 = 150.64 g/mol

= 150.64 × 575 × 10^-3

= 86.618 grams of CuCl2